OpenStudy (anonymous):

The equation passes through the point (5, -2) and is perpendicular to the line x - 2y - 4 = 0.

OpenStudy (anonymous):

First, solve explicitly for y: x-2y-4=0 2y=4-x y=2-(1/2)x m=-(1/2) The line passing through p(5, -2) is perpendicular, so the slop of new line will be the opposite reciprocal m=2 Using pt slope formula plug in your values; y-(-2)=2(x-5) solve for y y+2=2x-10 y=2x-12

OpenStudy (anonymous):

Refer to the Mathematica 9 solution and plot attached.